# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 5. A derived Milnor-Moore theorem
The initial $\mathrm{Comm}$-algebra ${\mathrm{H}}({\mathbb{Q}})$ in ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ (Lemma 2.24) has an underling $\mathrm{Comm}^{\mathrm{nu}}$-algebra used for the following lemma:
@@
A stable $\infty$-category ${\mathcal{C}}$ is ${\mathbb{Q}}$-linear if for any ${\mathrm{X}},{\mathrm{Y}}\in{\mathcal{C}}$, ${\mathrm{n}}>0$ the map
${\mathrm{n}}\bullet(-):\pi_{0}{\mathcal{C}}({\mathrm{X}},{\mathrm{Y}})\to\pi_{0}{\mathcal{C}}({\mathrm{X}},{\mathrm{Y}}),f\mapsto X\xrightarrow{}{\mathrm{X}}^{\oplus{\mathrm{n}}}\xrightarrow{f^{\oplus{\mathrm{n}}}}{\mathrm{Y}}^{\oplus{\mathrm{n}}}\to{\mathrm{Y}}$
@@
The next lemma is well-known to the experts and a proof is sketched in [11, Proposition 1.7.2.]. We offer an alternative proof for the reader’s convenience, which we find more conceptual and transparent.
@@
The $\mathrm{Comm}^{\mathrm{nu}}$-algebra ${\mathrm{H}}({\mathbb{Q}})[-1]$ in ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ is trivial.
@@
Let ${\mathrm{D}}$ be the rational chain complex concentrated in degree 00 and $-1$ whose non-trivial differential is the identity of ${\mathbb{Q}}.$ There is map $\kappa:{\mathrm{D}}\to{\mathbb{Q}}$ of rational chain complexes that is in degree zero the identity. Consider the pullback square of non-unital ${\mathbb{E}}_{\infty}$-algebras in rational chain complexes
(14) ${{\mathrm{P}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sym}^{\mathrm{nu}}({\mathrm{D}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${{\mathbb{Q}},}$
where ${\mathbb{Q}}$ carries the non-trivial commutative ${\mathbb{Q}}$-algebra structure underlying the unique commutative ${\mathbb{Q}}$-algebra structure, and the right vertical map is induced by $\kappa$ via the free non-unital ${\mathbb{E}}_{\infty}$-algebra $\mathrm{Sym}^{\mathrm{nu}}.$ The map $\phi$ is degree-wise surjective since $\kappa$ is degree-wise surjective. Therefore $\phi$ is a fibration of non-unital ${\mathbb{E}}_{\infty}$-algebras in rational chain complexes so that square (14) is a homotopy pullback square. Since ${\mathrm{D}}$ is acyclic and $\mathrm{Sym}^{\mathrm{nu}}$ preserves quasi-isomorphisms, $\mathrm{Sym}^{\mathrm{nu}}({\mathrm{D}})$ is acyclic. Hence ${\mathrm{P}}$ is the loops of ${\mathbb{Q}}$ in non-unital ${\mathbb{E}}_{\infty}$-algebras.
The unit ${\mathrm{D}}\to\mathrm{Sym}^{\mathrm{nu}}({\mathrm{D}})$ is an quasi-isomorphism since source and target are acyclic. Therefore the induced map $\lambda:{\mathbb{Q}}[-1]=\mathsf{fib}(\kappa)\to{\mathrm{P}}=\mathsf{fib}(\phi)$ on kernels is a quasi-isomorphism. Since ${\mathrm{P}}$ is a ${\mathbb{E}}_{\infty}$-algebra in rational chain complexes, the square of an element of ${\mathrm{P}}$ of degree -1 is zero. Thus the image of the generator under $\lambda$ squares to zero, so that $\lambda$ is a map of non-unital ${\mathbb{E}}_{\infty}$-algebras.
Thereore ${\mathbb{Q}}[-1]$ endowed with the trivial algebra structure is loops of ${\mathbb{Q}}$ in non-unital ${\mathbb{E}}_{\infty}$-algebras. The canonical symmetric monoidal functor
$\gamma:\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}}(\mathsf{Ch}_{{\mathbb{Q}}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})$
sends square (14) to a pullback square since the forgetful functor $\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ detects loops and the functor $\mathsf{Ch}_{{\mathbb{Q}}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ sends homotopy pullback squares to pullback squares. Hence $\gamma$ preserves trivial algebras, ${\mathrm{H}}({\mathbb{Q}})[-1]$ carries the trivial $\mathrm{Comm}^{\mathrm{nu}}$-algebra structure.
Notation 5.19.
@@
The next lemma is [3, Proposition 3.9.]:
@@
$\theta:\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$
@@
[…]
@@
[…]
@@
By Lemma 5.18 the $\mathrm{Comm}^{\mathrm{nu}}$-algebra ${\mathrm{H}}({\mathbb{Q}})[-1]$ is trivial. Hence the claim follows from the existence of the following commutative square provided by Corollary 5.21:
${\mathrm{Alg}_{{\mathrm{triv}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{{\mathrm{triv}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{triv}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nu}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}).}$
We obtain the following corollary, which is [11, Proposition 1.7.2.]:
@@
We are grateful to Gijs Heuts for communicating to us the following lemma, which also appears as [15, Lemma 13.8., 13.10.].
@@
$\alpha:{\mathrm{Y}}\to\lim_{{\mathrm{n}}\geq 1}\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{Y}}$
@@
$\alpha:{\mathrm{X}}\to\lim_{{\mathrm{n}}\geq 1}\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}_{\mathcal{O}}({\mathrm{X}})$
in ${\mathcal{C}}$ is an equivalence. By stability it is enough to see that the cofiber of $\alpha$ vanishes. The latter is the limit of the inverse system $\to{\mathrm{cofib}}(\alpha_{{\mathrm{n}}+1})\to{\mathrm{cofib}}(\alpha_{\mathrm{n}})\to...\to{\mathrm{cofib}}(\alpha_{1})$ of cofibers of the maps
$\alpha_{\mathrm{n}}:{\mathrm{X}}\to\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}_{\mathcal{O}}({\mathrm{X}})\simeq(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})\circ{\mathrm{X}}.$
Consequently, it is enough to see that for every ${\mathrm{n}}\geq 1$ the induced morphism ${\mathrm{cofib}}(\alpha_{{\mathrm{n}}+1})\to{\mathrm{cofib}}(\alpha_{\mathrm{n}})$ is the zero morphism.
Observe that $(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{1}\simeq{\mathbb{1}}$ and $\alpha_{\mathrm{n}}$ is the canonical morphism
${\mathrm{X}}\simeq(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{1}\otimes{\mathrm{X}}\to\coprod_{{\mathrm{k}}\geq 1}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{X}}^{\otimes{\mathrm{k}}}.$
Hence the cofiber of $\alpha_{\mathrm{n}}$ is equivalent to $\prod_{{\mathrm{k}}\geq 2}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{X}}^{\otimes{\mathrm{k}}}$ and the canonical morphism ${\mathrm{cofib}}(\alpha_{\mathrm{n}})\to{\mathrm{cofib}}(\alpha_{{\mathrm{n}}-1})$ identifies with the morphism
$\coprod_{{\mathrm{k}}\geq 2}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{X}}^{\otimes{\mathrm{k}}}\to\coprod_{{\mathrm{k}}\geq 2}({\mathcal{O}}_{\leq{\mathrm{n}}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{X}}^{\otimes{\mathrm{k}}}$
induced by the morphisms $(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\to({\mathcal{O}}_{\leq{\mathrm{n}}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}$ for ${\mathrm{k}}\geq 2.$
@@
    For every ${\mathrm{k}}>1$ the rational homology of $(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}$ is concentrated in degrees $\geq 1-{\mathrm{n}}$.
@@
    For every ${\mathrm{k}}>1$ the rational homology of $(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}$ is concentrated in degrees $\leq 1-{\mathrm{n}}$.
1: By assumption the rational homology of $(\tau_{1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\simeq({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\simeq{\mathcal{Q}}_{\mathrm{k}}$ is concentrated in degree zero. Therefore 1. follows by induction over ${\mathrm{n}}\geq 1$ from the fact that the rational homology of the fiber ${\mathrm{F}}$ of the map $(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}\to({\mathcal{O}}_{\leq{\mathrm{n}}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}$ is concentrated in degree $1-{\mathrm{n}}:$ the fiber ${\mathrm{F}}$ is the ${\mathrm{k}}$-th term of the cofree right ${\mathcal{Q}}\simeq{\mathcal{O}}^{\vee}$-comodule on the symmetric sequence concentrated in degree ${\mathrm{n}}$ with value ${\mathcal{O}}_{\mathrm{n}}.$ We set ${\mathcal{O}}^{\prime}:={\mathcal{O}}(-1)={\mathcal{Q}}^{\vee}(-1)$ so that ${\mathcal{O}}^{\prime}$ is an $\infty$-operad in ${\mathbb{Q}}$-vector spaces. Then we have that
${\mathrm{F}}\simeq({\mathcal{O}}^{\prime}_{\mathrm{n}}[1-{\mathrm{n}}]\otimes{\mathcal{Q}}^{\otimes{\mathrm{n}}})_{\mathrm{k}}\simeq\coprod_{{\mathrm{k}}_{1}+...+{\mathrm{k}}_{\mathrm{n}}={\mathrm{k}}}{\mathcal{O}}^{\prime}_{\mathrm{n}}[1-{\mathrm{n}}]\otimes\Sigma_{\mathrm{k}}\times_{(\Sigma_{{\mathrm{k}}_{1}}\times...\times\Sigma_{{\mathrm{k}}_{\mathrm{n}}})}({\mathcal{Q}}_{{\mathrm{k}}_{1}}\otimes...\otimes{\mathcal{Q}}_{{\mathrm{k}}_{\mathrm{n}}})$
has rational homology concentrated in degree $1-{\mathrm{n}}.$
2: Let $\tau_{>{\mathrm{n}}}({\mathcal{O}})$ be the fiber of the canonical map ${\mathcal{O}}\to\tau_{\mathrm{n}}({\mathcal{O}})$ of $\infty$-operads considered as a map of right ${\mathcal{O}}$-modules in ${\mathrm{sSeq}}(\mathrm{Sp}).$ Applying the exact functor $(-)\circ_{{\mathcal{O}}}{\mathrm{triv}}:{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}(\mathrm{Sp}))\to{\mathrm{sSeq}}(\mathrm{Sp})$ we get a fiber sequence ${\mathrm{W}}\to{\mathrm{triv}}\to\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}$ and so for every ${\mathrm{k}}>1$ a fiber sequence ${\mathrm{W}}_{\mathrm{k}}\to 0\to(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}}$. Hence ${\mathrm{W}}_{{\mathrm{k}}}\simeq\Omega((\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{\mathrm{k}})$. Consequently, 2. is equivalent to the condition that for every ${\mathrm{k}}>1$ the rational homology of ${\mathrm{W}}_{\mathrm{k}}=(\tau_{>{\mathrm{n}}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})_{{\mathrm{k}}}$ is concentrated in degrees $\leq-{\mathrm{n}}$.
By Remark 2.7 and Remark 5.24 we find that ${\mathrm{W}}_{\mathrm{k}}$ is the colimit of a filtered diagram
$\tau_{>{\mathrm{n}}}({\mathcal{O}})\simeq{\mathrm{D}}_{0}\to...\to{\mathrm{D}}_{\ell}\to...$

### v2

@@ 4. A derived Milnor-Moore theorem
Theorem 3.13 gives the following:
Theorem 4.17.
Let $\mathcal{C}$} be a stable presentably symmetric monoidal $\infty$-category such that all norm maps of objects in $\mathcal{C}$} with an action of a symmetric group are equivalences. For every Lie algebra $X\in{\mathcal{C}}$ the unit ${\mathrm{X}}\to{\mathrm{Prim}}{\mathcal{U}}(X)$ identifies with the completion morphism
$X\to\lim_{{n}\geq 1}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}X,$
where we identify $X$ via Corollary 4.9 with a group object in ${\mathrm{Lie}}(1)$-algebras.
Proof.
Let $X^{\prime}\in{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))$ be the group object in ${\mathrm{Lie}}(1)$-algebras in $\mathcal{C}$} corresponding to $X$ via the equivalence of Corollary 4.9.
The unit of Koszul duality
(4.3) $X^{\prime}\to{\mathrm{triv}}*^{{\mathrm{Lie}}(1)^{\vee}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}X^{\prime})$
is a morphism in ${\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))$ whose corresponding morphism of Lie-algebras in $\mathcal{C}$} is the unit $X\to{\mathrm{Prim}}{\mathcal{U}}(X)$ by construction of the adjunction ${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}.$
By Theorem 3.13 the morphism (4.3) is the completion morphism
$X^{\prime}\to\lim_{{n}\geq 1}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}X^{\prime}.$
The initial $\mathrm{Comm}$-algebra ${\mathrm{H}}({\mathbb{Q}})$ in ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ (Lemma 2.24) has an underling $\mathrm{Comm}^{\mathrm{nun}}$-algebra used for the following lemma:
@@
A stable $\infty$-category $\mathcal{C}$} is $\mathbb{Q}$}-linear if for any ${\mathrm{X}},{\mathrm{Y}}\in{\mathcal{C}}$, n>0{n}>0 the map
${n}\bullet(-):\pi_{0}{\mathcal{C}}({\mathrm{X}},{\mathrm{Y}})\to\pi_{0}{\mathcal{C}}({\mathrm{X}},{\mathrm{Y}}),f\mapsto X\xrightarrow{}{\mathrm{X}}^{\oplus{n}}\xrightarrow{f^{\oplus{n}}}{\mathrm{Y}}^{\oplus{n}}\to{\mathrm{Y}}$
@@
The next lemma is well-known to the experts and a proof is sketched in [18, Proposition 1.7.2.]. We offer an alternative proof for the reader’s convenience, which we find more conceptual and transparent.
@@
The $\mathrm{Comm}^{\mathrm{nun}}$-algebra ${\mathrm{H}}({\mathbb{Q}})[-1]$ in ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ is trivial.
@@
[…]
$\gamma:\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}(\mathsf{Ch}_{{\mathbb{Q}}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})$
sends square (4.4) to a pullback square since the forgetful functor $\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ detects loops and the functor $\mathsf{Ch}_{{\mathbb{Q}}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}$ sends homotopy pullback squares to pullback squares. Hence $\gamma$ preserves trivial algebras, ${\mathrm{H}}({\mathbb{Q}})[-1]$ carries the trivial $\mathrm{Comm}^{\mathrm{nun}}$-algebra structure.
Notation 4.22.
@@
The next lemma is [8, Proposition 3.9.]:
@@
$\theta:\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$
@@
$\{{\mathrm{H}}({\mathbb{Q}})[-1]\}\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$
@@
$\{{\mathrm{H}}({\mathbb{Q}})\}\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\xrightarrow{\Omega}\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}),$
@@
By Lemma 4.21 the $\mathrm{Comm}^{\mathrm{nun}}$-algebra ${\mathrm{H}}({\mathbb{Q}})[-1]$ is trivial. Hence the claim follows from the existence of the following commutative square provided by Corollary 4.24:
${\mathrm{Alg}_{{\mathrm{triv}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{{\mathrm{triv}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{triv}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})})\times\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{\mathrm{Comm}^{\mathrm{nun}}\boxtimes{\mathcal{O}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}).}$
We obtain the following corollary, which is [18, Proposition 1.7.2.]:
@@
We are grateful to Gijs Heuts for communicating to us the following lemma, which also appears as [31, Lemma 13.8., 13.10.].
@@
$\alpha:{\mathrm{Y}}\to\lim_{{n}\geq 1}\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{Y}}$
@@
$\alpha:{\mathrm{X}}\to\lim_{{n}\geq 1}\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}_{\mathcal{O}}({\mathrm{X}})$
in $\mathcal{C}$} is an equivalence. By stability it is enough to see that the cofiber of $\alpha$ vanishes. The latter is the limit of the inverse system $\to\mathrm{cofib}(\alpha_{{n}+1})\to\mathrm{cofib}(\alpha_{n})\to...\to\mathrm{cofib}(\alpha_{1})$ of cofibers of the maps
$\alpha_{n}:{\mathrm{X}}\to\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}_{\mathcal{O}}({\mathrm{X}})\simeq(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})\circ{\mathrm{X}}.$
Consequently, it is enough to see that for every ${n}\geq 1$ the induced morphism $\mathrm{cofib}(\alpha_{{n}+1})\to\mathrm{cofib}(\alpha_{n})$ is the zero morphism.
Observe that $(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{1}\simeq{\mathbb{1}}$ and $\alpha_{n}$ is the canonical morphism
${\mathrm{X}}\simeq(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{1}\otimes{\mathrm{X}}\to\coprod_{{k}\geq 1}(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\otimes_{\Sigma_{k}}{\mathrm{X}}^{\otimes{k}}.$
Hence the cofiber of $\alpha_{n}$ is equivalent to $\prod_{{k}\geq 2}(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\otimes_{\Sigma_{k}}{\mathrm{X}}^{\otimes{k}}$ and the canonical morphism $\mathrm{cofib}(\alpha_{n})\to\mathrm{cofib}(\alpha_{{n}-1})$ identifies with the morphism
$\coprod_{{k}\geq 2}(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\otimes_{\Sigma_{k}}{\mathrm{X}}^{\otimes{k}}\to\coprod_{{k}\geq 2}({\mathcal{O}}_{\leq{n}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\otimes_{\Sigma_{k}}{\mathrm{X}}^{\otimes{k}}$
induced by the morphisms $(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\to({\mathcal{O}}_{\leq{n}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{k}$ for ${k}\geq 2.$
@@
    For every k>1{k}>1 the rational homology of $(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}$ is concentrated in degrees $\geq 1-{n}$.
@@
    For every k>1{k}>1 the rational homology of $(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}$ is concentrated in degrees $\leq 1-{n}$.
1: By assumption the rational homology of $(\tau_{1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\simeq({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\simeq{\mathcal{Q}}_{k}$ is concentrated in degree zero. Therefore 1. follows by induction over ${n}\geq 1$ from the fact that the rational homology of the fiber $\mathrm{F}$} of the map $(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}\to({\mathcal{O}}_{\leq{n}-1}\circ_{\mathcal{O}}{\mathrm{triv}})_{k}$ is concentrated in degree 1n:1-{n}: the fiber $\mathrm{F}$} is the k{k}-th term of the cofree right ${\mathcal{Q}}\simeq{\mathcal{O}}^{\vee}$-comodule on the symmetric sequence concentrated in degree n{n} with value ${\mathcal{O}}_{n}.$ We set ${\mathcal{O}}^{\prime}:={\mathcal{O}}(-1)={\mathcal{Q}}^{\vee}(-1)$ so that $\mathcal{O}$$^{\prime}$ is an $\infty$-operad in $\mathbb{Q}$}-vector spaces. Then we have that
${\mathrm{F}}\simeq({\mathcal{O}}^{\prime}_{n}[1-{n}]\otimes{\mathcal{Q}}^{\otimes{n}})_{k}\simeq\coprod_{{k}_{1}+...+{k}_{n}={k}}{\mathcal{O}}^{\prime}_{n}[1-{n}]\otimes\Sigma_{k}\times_{(\Sigma_{{k}_{1}}\times...\times\Sigma_{{k}_{n}})}({\mathcal{Q}}_{{k}_{1}}\otimes...\otimes{\mathcal{Q}}_{{k}_{n}})$
has rational homology concentrated in degree 1n.1-{n}.
2: Let $\tau_{>{n}}({\mathcal{O}})$ be the fiber of the canonical map ${\mathcal{O}}\to\tau_{n}({\mathcal{O}})$ of $\infty$-operads considered as a map of right $\mathcal{O}$}-modules in ${\mathrm{sSeq}}({\mathrm{Sp}}).$ Applying the exact functor $(-)\circ_{{\mathcal{O}}}{\mathrm{triv}}:{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathrm{Sp}}))\to{\mathrm{sSeq}}({\mathrm{Sp}})$ we get a fiber sequence ${\mathrm{W}}\to{\mathrm{triv}}\to\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}$ and so for every k>1{k}>1 a fiber sequence ${\mathrm{W}}_{k}\to 0\to(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k}$. Hence ${\mathrm{W}}_{{k}}\simeq\Omega((\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}})_{k})$. Consequently, 2. is equivalent to the condition that for every k>1{k}>1 the rational homology of ${\mathrm{W}}_{k}=(\tau_{>{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})_{{k}}$ is concentrated in degrees $\leq-{n}$.
By Remark 2.7 and Remark 4.27 we find that ${\mathrm{W}}_{k}$ is the colimit of a filtered diagram
$\tau_{>{n}}({\mathcal{O}})\simeq{\mathrm{D}}_{0}\to...\to{\mathrm{D}}_{\ell}\to$...